# Size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction method for the supramolecular approach

### v1

@@ IV.1 Size consistency of sc-HSB-QSCI
Table 1: Sc-HSB-QSCI, RHF, and full-CI energies for the 8H (dimer) and 4H (monomer) clusters, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants included in the subspace Hamiltonian diagonalization.
$K$ EE(dimer) EE(monomer 1) EE(monomer 2)
1 -3.652202 (98) -1.816720 (8) -1.835483 (8)
2 -3.862635 (124) -1.933870 (9) -1.928764 (10)
3 -3.862635 (124) -1.933870 (9) -1.928764 (10)
4 -3.862664 (134) -1.933900 (10) -1.928764 (10)
5 -3.862664 (134) -1.933900 (10) -1.928764 (10)
RHF -3.553541 (1) -1.776770 (1) -1.776770 (1)
full-CI -3.878863 (178) -1.939432 (12) -1.939432 (12)
@@
Here, we numerically demonstrate that the sc-HSB-QSCI method satisfies the size consistency condition, using the 4H cluster as the monomer. The square 4H cluster is a well-known strongly correlated system [45], with a ground-state wave function exhibiting significant multi-configurational character. Using the STO-3G basis set [35], the full-CI dimensions are 12 and 178 for the monomer (4H cluster in the $D_{2h}$ point group) and the dimer (8H cluster in the $C_{2v}$ point group), respectively. Owing to the small system size, sc-HSB-QSCI calculations were performed with 10 shots for each $k$. The results are summarized in Table 1. The computed $E_{\text{int}}$ values are less than $1\times 10^{-7}$ Hartree for all $K$. Tne small deviations of $E_{\text{int}}$ from zero arize from rounding errors in the one- and two-electron integrals. These results clearly demonstrate the size consistent nature of sc-HSB-QSCI.
@@
Next, we numerically investigated size consistency in the FH dimer and FH–H2O systems using both sc-HSB-QSCI and the original form of HSB-QSCI (org-HSB-QSCI). As discussed above, org-HSB-QSCI is not size consistent, and the interaction energy can be nonzero even when the intermolecular distance is sufficiently large. In this study, we used geometries with an intermolecular distance of 100 Å. The calculated interaction energies are summarized in Table 2, and total energies and the number of Slater determinants in the subspace are provided in Tables S3 and S4 in the Supporting Information. Sc-HSB-QSCI yielded an interaction energy $E_{\text{int}}$ = 0.000 kcal mol-1 for all $K$, whereas org-HSB-QSCI gave nonzero interaction energies. The error in $E_{\text{int}}$ is slightly larger for FH–H2O than for the FH dimer, suggesting that the violation of size consistency can increase as the number of active orbitals grows. In all systems under study, the number of sampled charge-transfer excited Slater determinants was zero.
IV.2 Intermolecular interaction energies of hydrogen-bonded FH dimer and FH–H2O
Lastly, we examine the intermolecular interaction energy calculations of the hydrogen-bonded FH dimer and FH–H2O. Hydrogen bonding is one of the most important intermolecular interactions, particularly in biomolecules. Here, we compare interaction energies calculated using the supramolecular approach with sc-HSB-QSCI and org-HSB-QSCI, as well as the dimer approach with org-HSB-QSCI. The numerical simulation results, together with RHF and CAS-CI interaction energies, are summarized in Table 3, and the total energies are listed in Tables S5 and S6 in the Supporting Information. Sc-HSB-QSCI yields larger interaction energies in absolute value and smaller errors relative to the CAS-CI values compared to org-HSB-QSCI. Org-HSB-QSCI using the dimer approach gave interaction energies comparable to those obtained with sc-HSB-QSCI, although its computational cost is higher than that of the supramolecular approach. At $K=10$, sc-HSB-QSCI gave interaction energies with deviations of 0.012 and 0.038 kcal mol-1 from the CAS-CI values for the FH dimer and FH–H2O, respectively. The number of Slater determinants included in the subspace Hamiltonian diagonalization for the dimer is 458 and 4150 for the FH dimer and FH–H2O, respectively, which should be compared to 2025 and 27225, respectively, at the CAS-CI level of theory.
@@ IV.2 Intermolecular interaction energies of hydrogen-bonded FH dimer and FH–H2O
Since our size-consistent implementation is applicable to other selected CI frameworks, we applied our proposed scheme to heat-bath CI (HCI) [46]. Details of the numerical simulations are provided in subsection C of the Supporting Information. The size consistent HCI (sc-HCI) yielded intermolecular interaction energies slightly closer to the CAS-CI values than those obtained with sc-HSB-QSCI, indicating that sc-HCI achieves a more compact wave function expansion for a given level of accuracy. As discussed in our previous work [10], HSB-QSCI may sample Slater determinants that are important for excited states rather than the ground state. Therefore, screening out unimportant Slater determinants from the subspace is essential for obtaining a more compact wave function expansion in sc-HSB-QSCI, which remains one of the most significant challenges in QSCI. Nevertheless, the approach proposed in this study significantly improves the applicability of selected CI methods to intermolecular interaction energy calculations.
V Conclusion
In summary, we have developed a size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction (sc-HSB-QSCI) for intermolecular interaction energy calculations based on the supramolecular approach. The size consistency of the proposed method was confirmed through numerical simulations of the 8H cluster, FH dimer, and FH–H2O systems with an inter-monomer distance of 100 Å, all yielding zero interaction energies. Interaction energies of intermolecular hydrogen bonds in the FH dimer and FH–H2O were also calculated with sc-HSB-QSCI, and the results agreed with the CAS-CI interaction energies in the same active space with errors of only 0.012 and 0.038 kcal mol-1 for the FH dimer and FH–H2O, respectively. During the sampling procedure, intra-monomer excited configurations are predominantly sampled in the early stage of real-time evolution, while the probability of sampling charge-transfer excited configurations increases at later stages. This trend indicates that extended real-time evolution is essential for accurately computing intermolecular interaction energies with sc-HSB-QSCI.
Apart from the supramolecular approach, size consistency is also crucial in other contexts, such as the fragment molecular orbital method [47, 48] and the ONIOM (our own N-layer integrated molecular orbitals and molecular mechanics) method [49, 50]. Applying the sc-HSB-QSCI method to these problems remains an important avenue for future research.
VI Acknowledgments

### v2

@@ V.1 Size consistency of sc-HSB-QSCI
Table 1: Sc-HSB-QSCI, RHF, and full-CI total energies, errors in the total energy with respect to the full-CI value ($\Delta E_{\text{sc-HSB-QSCI}-\text{full-CI}}$/Hartree) for the 8H (dimer) and 4H (monomer) clusters, in units of Hartree, and intermolecular interaction energy ($E_{int}$) in units of kcal mol-1. The numbers in parenthesis of the total energy section indicate the number of Slater determinants included in the subspace Hamiltonian diagonalization.
$K$ Total energy/Hartree $\Delta E_{\text{sc-HSB-QSCI}-\text{full-CI}}$/Hartree $E_{int}$/kcal mol-1
Dimer Monomer 1 Monomer 2 Dimer Monomer 1 Monomer 2
1 -3.652202 (98) -1.816720 (8) -1.835483 (8) 0.226661 0.122712 0.103949 0.0000
2 -3.862635 (124) -1.933870 (9) -1.928764 (10) 0.016229 0.005561 0.010667 0.0000
3 -3.862635 (124) -1.933870 (9) -1.928764 (10) 0.016229 0.005561 0.010667 0.0000
4 -3.862664 (134) -1.933900 (10) -1.928764 (10) 0.016199 0.005532 0.010667 0.0000
5 -3.862664 (134) -1.933900 (10) -1.928764 (10) 0.016199 0.005532 0.010667 0.0000
RHF -3.553541 (1) -1.776770 (1) -1.776770 (1) 0.0000
full-CI -3.878863 (178) -1.939432 (12) -1.939432 (12) 0.0000
@@
Here, we demonstrate that the sc-HSB-QSCI method satisfies the size consistency condition, in the 8H cluster using the 4H cluster as the monomer. The square 4H cluster is a well-known strongly correlated system [Paldus-1993], with a ground-state wave function exhibiting significant multi-configurational character. Using the STO-3G basis set [STO-3G], the full-CI dimensions are 12 and 178 for the monomer (4H cluster in the $D_{2h}$ point group) and the dimer (8H cluster in the $C_{2v}$ point group), respectively. Owing to the small system size, sc-HSB-QSCI calculations were performed with 10 shots for each $k$. The results are summarized in Table 1. The computed $E_{\text{int}}$ values are less than $1\times 10^{-7}$ Hartree for all $K$. The small deviations of $E_{\text{int}}$ from zero arize from rounding errors in the one- and two-electron integrals. These results clearly demonstrate the size consistent nature of sc-HSB-QSCI. Note that the 8H cluster consists of two identical monomers (4H clusters), yet the sc-HSB-QSCI energies in Table 1 do not yield identical energies for the two monomers. This discrepancy arises because the sampling process in sc-HSB-QSCI is inherently stochastic, and symmetrically equivalent molecules can exhibit different energies if the sets of sampled Slater determinants differ. Although point group symmetry is not considered in this work, it is possible to leverage molecular point group symmetry to ensure that two equivalent molecules have the same energies.
Importantly, the present implementation consistently satisfies the size consistency condition, regardless of the system size or the number of monomers, because the electronic configurations required for size consistency are automatically included in the subspace. To demonstrate this, we performed numerical simulations for the 12H cluster as the trimer and evaluated the size consistency error by calculating the following quantities.
$\Delta E_{123}=E(123)-E(1)-E(2)-E(3)$ (3)
$\Delta E_{12}=E(12)-E(1)-E(2)$ (4)
$\Delta E_{13}=E(13)-E(1)-E(3)$ (5)
$\Delta E_{23}=E(23)-E(2)-E(3)$ (6)
Here, $E(1)$, $E(2)$, and $E(3)$ denote the total energies of monomers 1, 2, and 3, respectively, $E(12)$, $E(13)$, and $E(23)$ represent the energies of the dimers composed of monomers 1 and 2, 1 and 3, and 2 and 3, respectively, and $E(123)$ is the energy of the trimer. Table 2 summarizes the simulation results, which yield zero intermolecular interaction energies for all systems.
Table 2: Intermolecular interaction energies for the 12H cluster consisting of three 4H clusters at inter-monomer distances of 100 Å  in units of kcal mol-1.
$K$ $\Delta E_{123}$ $\Delta E_{12}$ $\Delta E_{13}$ $\Delta E_{23}$
1 0.0000 0.0000 0.0000 0.0000
2 0.0000 0.0000 0.0000 0.0000
3 0.0000 0.0000 0.0000 0.0000
4 0.0000 0.0000 0.0000 0.0000
5 0.0000 0.0000 0.0000 0.0000
10 0.0000 0.0000 0.0000 0.0000
@@
Next, we numerically investigated size consistency in the FH dimer and FH–H2O systems using both sc-HSB-QSCI and the original form of HSB-QSCI (org-HSB-QSCI). As discussed above, org-HSB-QSCI is not size consistent, and the interaction energy can be nonzero even when the intermolecular distance is sufficiently large. In this study, we used geometries with an intermolecular distance of 100 Å. The calculated interaction energies are summarized in Table 3, and total energies and the number of Slater determinants in the subspace are provided in Tables S3 and S4 in the Supporting Information. Sc-HSB-QSCI yielded an interaction energy $E_{\text{int}}$ = 0.000 kcal mol-1 for all $K$, whereas org-HSB-QSCI gave nonzero interaction energies. The error in $E_{\text{int}}$ is slightly larger for FH–H2O than for the FH dimer, suggesting that the violation of size consistency can increase as the number of active orbitals grows. It should also be noted that the energy error arising from violation of size consistency in org-HSB-QSCI is less than 0.1 kcal mol-1 in Table 3, but this is because a relatively large number of shots was used. Table 4 summarizes the dependence of the intermolecular interaction energies on the number of shots for the FH dimer and FH–H2O at an inter-monomer distance of 100 Å. It is clear that violation of size consistency become more pronounced when a smaller number of shots is employed. There results also suggest that violation of size consistency becomes more significant as the active space increases. In all systems under study, the number of sampled charge-transfer excited Slater determinants was zero.
Table 4: Shot number dependence on intermolecular interaction energies for the FH dimer and FH–H2O at an inter-monomer distance of 100 Å, calculated using sc-HSB-QSCI and org-HSB-QSCI at $K=5$, in units of kcal mol-1.
Molecule #shots sc-HSB-QSCI org-HSB-QSCI
FH dimer 100 0.000 0.540
200 0.000 0.447
300 0.000 -0.059
400 0.000 0.085
500 0.000 -0.221
1000 0.000 0.231
10000 0.000 0.007
FH–H2O 100 0.000 0.710
200 0.000 0.471
300 0.000 -0.631
400 0.000 0.344
500 0.000 0.193
1000 0.000 -0.143
10000 0.000 0.031
V.2 Intermolecular interaction energies of hydrogen-bonded FH dimer and FH–H2O
Next, we examine the intermolecular interaction energy calculations of the hydrogen-bonded FH dimer and FH–H2O. Hydrogen bonding is one of the most important intermolecular interactions, particularly in biomolecules. Here, we compare interaction energies calculated using the supramolecular approach with sc-HSB-QSCI and org-HSB-QSCI, as well as the dimer approach with org-HSB-QSCI. The numerical simulation results, together with RHF and CAS-CI interaction energies, are summarized in Table 5, and the total energies are listed in Tables S5 and S6 in the Supporting Information. Sc-HSB-QSCI yields larger interaction energies in absolute value and smaller errors relative to the CAS-CI values compared to org-HSB-QSCI. Org-HSB-QSCI using the dimer approach gave interaction energies comparable to those obtained with sc-HSB-QSCI, although its computational cost is higher than that of the supramolecular approach. At $K=10$, sc-HSB-QSCI gave interaction energies with deviations of 0.012 and 0.038 kcal mol-1 from the CAS-CI values for the FH dimer and FH–H2O, respectively. The number of Slater determinants included in the subspace Hamiltonian diagonalization for the dimer is 458 and 4150 for the FH dimer and FH–H2O, respectively, which should be compared to 2025 and 27225, respectively, at the CAS-CI level of theory.
@@ V.2 Intermolecular interaction energies of hydrogen-bonded FH dimer and FH–H2O
In the hydrogen-bonded FH dimer and FH–H2O systems, one unoccupied molecular orbital is delocalized over the dimer (See the 10th and 11th LMOs of the FH dimer and FH–H2O, respectively, in Figure S2 in the Supporting Information). The overlaps between the molecular orbitals of the monomer and dimer, $|\langle\varphi_{i;\text{Monomer}}|\varphi_{j;\text{Dimer}}\rangle|$, are calculated to be 0.959 and 0.840 for the FH dimer and FH–H2O, respectively. Such molecular orbital delocalization indicates that the monomer and dimer do not share the same Hilbert space as the active space, which can influence the calculated intermolecular interaction energy. Generating nearly identical LMOs for both the monomer and dimer when the two molecules are in close proximity remains an unresolved issue. However, as discussed in the Theory section, this problem may be mitigated by designing a cost function for molecular orbital localization that includes a penalty when molecular orbitals are delocalized over the dimer. In contrast, in the absence of intermolecular interactions, such orbital delocalization does not occur, and the intermolecular interaction energy becomes zero.
@@
Since our size-consistent implementation is applicable to other selected CI frameworks, we applied our proposed scheme to HCI [Holmes-2016]. Details of the numerical simulations are provided in subsection C of the Supporting Information. The size consistent HCI (sc-HCI) yielded intermolecular interaction energies slightly closer to the CAS-CI values than those obtained with sc-HSB-QSCI, indicating that sc-HCI achieves a more compact wave function expansion for a given level of accuracy. As discussed in our previous work [Sugisaki-2025], HSB-QSCI may sample Slater determinants that are important for excited states rather than the ground state. Therefore, screening out unimportant Slater determinants from the subspace is essential for obtaining a more compact wave function expansion in sc-HSB-QSCI, which remains one of the most significant challenges in QSCI. Nevertheless, the approach proposed in this study significantly improves the applicability of selected CI methods to intermolecular interaction energy calculations.
VI Conclusion
In summary, we have developed a size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction (sc-HSB-QSCI) for intermolecular interaction energy calculations based on the supramolecular approach. The size consistency of the proposed method was confirmed through numerical simulations of the 8H and 12H clusters, FH dimer, and FH–H2O systems with an inter-monomer distance of 100 Å, all yielding zero intermolecular interaction energies. Interaction energies of intermolecular hydrogen bonds in the FH dimer and FH–H2O were also calculated with sc-HSB-QSCI, and the results agreed with the CAS-CI interaction energies in the same active space with errors of only 0.012 and 0.038 kcal mol-1 for the FH dimer and FH–H2O, respectively. During the sampling procedure, intra-monomer excited configurations are predominantly sampled in the early stage of real-time evolution, while the probability of sampling charge-transfer excited configurations increases at later stages. This trend indicates that extended real-time evolution is essential for accurately computing intermolecular interaction energies with sc-HSB-QSCI.
Apart from the supramolecular approach, size consistency is also crucial in other contexts, such as the fragment molecular orbital method [Kitaura-1999, Mochizuki-2022] and the ONIOM (our own N-layer integrated molecular orbitals and molecular mechanics) method [Svensson-1996, Chung-2015]. Applying the sc-HSB-QSCI method to these problems remains an important avenue for future research.
VII Acknowledgments
@@
VIII Supporting Information
VIII.1 Cartesian coordinates
Table S1: Cartesian coordinates of the PBE0/aug-$c$-pVDZ optimized geometry for the hydrogen-bonded FH dimer, given in units of Å.
Atom X Y Z
F -0.355909 0.404037 0.199053
H 0.112061 -0.309544 -0.156637
F -3.008949 -0.078707 -0.096402
H -2.130203 0.184214 0.053986
Table S2: Cartesian coordinates of the PBE0/aug-$c$-pVDZ optimized geometry for the hydrogen-bonded FH–H2O, given in units of Å.
Atom X Y Z
O -2.451734 0.852831 -0.885154
H -1.585104 0.709253 -0.491503
H -2.774601 1.674026 -0.500774
F -4.071275 -1.154030 -0.438062
H -3.491540 -0.434027 -0.626687
VIII.2 Relationship between the number of quantum circuit repetitions and size consistency
Here, we demonstrate that the number of quantum circuit repetitions must increase exponentially with system size in order to approximately satisfy the size consistency condition in QSCI. For clarity, our discussion focuses on the original QSCI and assumes access to a state preparation oracle capable of generating the CAS-CI wave function.
$Prep|0\rangle^{\otimes n}=|\Psi\rangle=\sum_{j}c_{j}|\psi_{j}\rangle$
Here, $|\Psi\rangle$ denotes the CAS-CI wave function of the target electronic state, and $\{|\psi_{j}\rangle\}$ represents the set of Slater determinants. In addition, we assume an ideal measurement scenario in which each Slater determinant $|\psi_{j}\rangle$ is sampled once for every $1/|c_{j}|^{2}$ measurements performed.
Let us consider a dimer composed of two monomers with no intermolecular interactions. In this case, the wave function of the dimer can be expressed as the direct product of the wave functions of the individual monomers, as shown below.
$|\Psi_{\text{dimer}}\rangle=\sum_{jk}c_{j}c_{k}|\psi_{j}\rangle\otimes|\psi_{k}\rangle$
Based on this, sampling the Slater determinant represented as the direct product $|\psi_{j}\rangle\otimes|\psi_{j}\rangle$ requires executing the quantum circuit $1/|c_{j}|^{4}$ times.
VIII.3 Size-consistent implementation for heat-bath CI
Figure S1: Number of dimer Slater determinants included in the subspace Hamiltonian versus intermolecular interaction energy obtained from sc-HCI and sc-HSB-QSCI in (a) FH dimer and in (b) FH–H2O.
As discussed in the Theory section, the size-consistent implementation proposed in this study is applicable to any selected CI framework. Here, we examine its application to heat-bath CI (sc-HCI). In sc-HCI, we first perform HCI calculations for the dimer, and the resulting Slater determinants are used to construct the subspaces for both the monomers and the dimer. In this work, HCI was carried out with threshold values $\varepsilon$ ranging from $1.0\times 10^{-4}$ to $5.0\times 10^{-3}$. Figure S1 plots the number of dimer Slater determinants against the intermolecular interaction energies calculated using both sc-HCI as well as sc-HSB-QSCI. Notably, sc-HCI yielded intermolecular interaction energies that are closer to the CAS-CI values than those obtained with sc-HSB-QSCI.